OpenStudy (macgirl234):

Given the function f(x) = 3(x+2) − 4, solve for the inverse function when x = 2. −4 0 4 8

OpenStudy (3mar):

If I may help?

OpenStudy (macgirl234):

Please do

OpenStudy (3mar):

Thank you.

OpenStudy (3mar):

Can you find out the inverse o this function firstly?

OpenStudy (macgirl234):

I know that the inverse can be found but I am unsure how to

OpenStudy (helper101):

can someone help me

OpenStudy (3mar):

Can you share what you got and I check it for you?

OpenStudy (helper101):

Can anyone help me with a doc

OpenStudy (macgirl234):

I didn't get an answer because I'm unsure how to solve for the inverse of f(x)

OpenStudy (3mar):

Of course, which is?

OpenStudy (3mar):

@macgirl234 Do you follow?

OpenStudy (3mar):

Can we proceed here? http://openstudy.com/study#/updates/581ba2cce4b03ac563220145

OpenStudy (macgirl234):

sure.

OpenStudy (3mar):

Sorry for misunderstanding!

OpenStudy (macgirl234):

no problem!

OpenStudy (macgirl234):

Could you give me the formula or steps to follow to solve for the inverse of f(x) then I can attempt to apply it to the problem?

OpenStudy (3mar):

Welcome back! ♥ ♥ ♥

OpenStudy (3mar):

Of course. With my pleasure!

OpenStudy (3mar):

Here are the steps: 1. If you are given a function y(x), first off you solve it for y, which means separate y in one side and the other terms in the other side. 2. solve for x, also separate x in one side and the other terms in the other side including y. 3. swap x and y.......you get the invrse. \[\huge hint\] the inverse of a function is the reflection of the original function about the line (y=x). That is for you to simply imagine it and makes sense for you!!

OpenStudy (macgirl234):

So, I would solve for y, solve for x, then flip them?

OpenStudy (3mar):

For our case: f(x) = 3(x+2) − 4 \[y=3(x+2)−4\] solve for x: \[y+4=3(x+2)−4+4\] \[y+4=3(x+2)\] \[y+4=3x+6\] \[y+4-6=3x+6-6\] \[y+4-6=3x\] \[\frac{ y-2 }{ 3 }= \frac{ 3x }{ 3 }\] \[\huge x=\frac{ y-2 }{ 3 }\] now swap x and y:::: \[\huge y=\frac{ x-2 }{ 3 }\] This is the inverse or the main function you are given!

OpenStudy (3mar):

@macgirl234 Satisfied?

OpenStudy (macgirl234):

Ahhhhhh!!! and then if x=2, that means that y= 2-2/3 y=0/3 y=0 Correct?

OpenStudy (3mar):

Excellent!

OpenStudy (macgirl234):

Woohoo!!! Thanks so much!

OpenStudy (3mar):

Thank you for learning!

OpenStudy (3mar):

Any more questions?

OpenStudy (macgirl234):

Not with this question but I do have other questions for different problems.

OpenStudy (macgirl234):

I wouldn't want to take up all of your time though.

OpenStudy (3mar):

At your service!

OpenStudy (macgirl234):

Are you sure?

OpenStudy (3mar):

Let me surprise you!